# anchor decoding AGC automatic gain control

### v1

@@ Abstract
Determining the exact decoding error probability of linear block codes is an interesting problem. For binary BCH codes, McEliece derived methods to estimate the error probability of a simple bounded distance decoding (BDD) for BCH codes. However, BDD falls short in many applications. In this work, we consider error-and-erasure decoding and its variants that improve upon BDD. We derive closed-form expressions for their error probabilities and validate them through simulations. Then, we illustrate their use in assessing concatenated coding schemes.
@@ I Introduction
One appealing property of BDD is that its decoding error probability can be determined analytically [8]: given the number of input errors, the distribution of residual errors after decoding is known. In contrast, no such results exist yet for EaE decoding. In this paper, we derive a closed-form expression for the considered EaED. Specifically, we aim to determine the decoding transition probability (DTP) P(R=r|U=u,E=e){P(R\!=\!r|U\!=\!u,E\!=\!e)}, where $R$ is the number of residual errors after decoding, $U$ the number of input errors, and $E$ the number of input erasures. With the DTP, the post-decoding bit error rate (BER) and frame error rate (FER) for various memoryless channels can be computed. The DTP is validated through simulations and applied in a concatenated coding scheme.
@@ II Preliminaries
Let $\mathcal{C}[n,k,d_{\textup{des}}]$ be a primitive, narrow-sense, binary BCH code, where $n$ is the block length, $k$ the dimension, and $d_{\textup{des}}$ the designed distance, a lower bound on the minimum distance $d_{\textnormal{min}}$. The code rate is given by $R_{c}=k/n$. Let $t$ denote the number of correctable errors. The parameters are related by $n=2^{b}-1$, $k\geq n-bt$, and $2t+1=d_{\textup{des}}$. Let $A_{w}$ denote the number of codewords of weight $w$ in $\mathcal{C}$. Binary primitive BCH codes have approximately binomial weight distributions given by $A_{w}\approx 2^{-bt}\binom{n}{w}$ for $d_{\mathrm{des}}\leq w\leq n-d_{\mathrm{des}}$, with $A_{0}=A_{n}=1$, and $A_{w}=0$ elsewhere [1]. The exact weight enumerators of some short BCH codes are listed in [9].
@@
We use $\bm{x}\in\mathcal{C}$ to denote the transmitted codeword and $\bm{y}$ to denote the received word, whose alphabet is determined by the channel. At the output of a BSC, let $\mathcal{S}_{t}(\bm{c})=\{\bm{y}\in\mathbb{F}_{2}^{n}:d(\bm{y},\bm{c})\leq t\}$ denote the Hamming sphere of radius $t$ for $\bm{c}\in\mathcal{C}$. Then
$\textsf{BDD}\left(\bm{y}\right)=\begin{cases}\bm{c},&\text{if }\bm{y}\in\mathcal{S}_{t}(\bm{c}),\\ \bm{y}\,(\mathsf{fail}),&\text{otherwise}.\end{cases}$
Note that the decoding outcome falls into one of three categories: (i) decoding success, if $\bm{c}=\bm{x}$; (ii) decoding failure if no codeword is found; or (iii) miscorrection (also called a decoding error as in [8]), if $\bm{c}\neq\bm{x}$.
@@
Note that for a decoding failure, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(R\neq u|u)=0$, since $\textsf{BDD}\left(\bm{y}\right)=\bm{y}$ in this case, and for a decoding success, $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(R\neq 0|u)=0$. Furthermore, we denote by $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(u)$, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u)$, and $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(u)$ the conditional probabilities of success, failure, and miscorrection given $u$ input errors, respectively.
@@ III Analysis of Bounded Distance Decoding
We define $a=|\{i\in[n]:c_{i}=1,y_{i}=0\}|$} as the number of correct bits in $\bm{y}$ that are (wrongly) flipped by BDD, and $b=|\{i\in[n]:c_{i}=0,y_{i}=1\}|$} as the number of erroneous positions in $\bm{y}$ that are (correctly) flipped by BDD. Hence, it holds that u+ab=r{u+a-b=r} and $a+b=d\left(\bm{y},\bm{c}\right)\leq t$. For fixed $a$ and $b$, there are $\binom{r}{a}\binom{n-r}{b}=\binom{r}{a}\binom{n-r}{u+a-r}$ distinct patterns for $\bm{y}$. Additionally, the number of codewords $\bm{c}\in\mathcal{C}$ with $\textnormal{wt}\left(\bm{c}\right)=r,$ is given by the weight enumerator $A_{r}$. Thus, we obtain
@@ IV-B Decoding Transition Probability
$P_{\textnormal{{EaED}}}^{d}(r|u,e)=\textstyle\sum_{e_{1}=0}^{e}P_{\textnormal{{EaED}}}^{d}(r|u,e,e_{1})\cdot P(e_{1}|u,e),$
@@
Using the $[255,239,5]$ BCH code as an example, Tab. I and Tab. II list the conditional probabilities $P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(u,e)$ and $P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(u,e)$, respectively, which are calculated using (2) and marginalization over $U$. Note that if $e=0$, Theorem 2 reverts to Theorem 1. Accordingly, the first column of both tables corresponds to BDD. The miscorrection rate of BDD is high and is close to the upper bound $\frac{1}{t!}$ derived in [8]. Compared with BDD, EaED can decode beyond the minimum distance. When $2u+e$ slightly exceeds $d_{\textnormal{min}}-1$, the decoding success rate of EaED is still high. However, this comes at the expense of an increased probability of miscorrection for larger values of $2u+e$. The post-decoding BERs of EaED and BDD are compared in Fig. 2 with the threshold $\mathsf{T}$ optimized to minimize the BER of EaED at $9$ dB. We see that the simulation and computed values closely agree. Additionally, as expected, EaED improves upon BDD in terms of performance.
@@ V-C Improving one-shot decoding of BCH code
As shown in Fig. 2, EaEDa outperforms both EaED and BDD. Next, we examine the miscorrection rate. For EaED, most bit errors are due to miscorrections, whereas BDD exhibits a smaller fraction. This observation aligns with Tab. I and Tab. II, reflecting that EaED has a stronger error correction capability but is more prone to miscorrection. By contrast, EaEDa substantially reduces miscorrections through the use of anchor bits. Thus, EaEDa improves decoding performance while maintaining a low miscorrection rate, keeping the undetectable error rate negligible.
@@ Appendix A EaED Calculation
Finally, we consider the cases in $\mathcal{M}$. First, note that they can only cause a miscorrection or a failure since $\bm{y}^{(1)},\bm{y}^{(2)}\notin\mathcal{S}_{t}({\bm{0}})$. In the following, we make a significant simplification by assuming that the outcomes of $\textsf{BDD}\left(\bm{y}^{(1)}\right)$ and $\textsf{BDD}\left(\bm{y}^{(2)}\right)$, whether miscorrection or failure, are independent of each other. In principle, this assumption is false as $d(\bm{y}^{(1)},\bm{y}^{(2)})=e$. However, as the positions of the erasures are random, we observe that fixing the decoding result of one BDD does not significantly influence the other one. Then, we obtain

### v2

@@ Abstract
Determining the exact decoding error probability of linear block codes is an interesting problem. For binary BCH codes, McEliece derived methods to estimate the error probability of a simple bounded distance decoding (BDD) for BCH codes. However, BDD falls short in many applications. In this work, we consider error-and-erasure decoding and its improved variants. We derive closed-form expressions for their error probabilities and validate them through simulations. Then, we illustrate their use in assessing concatenated coding schemes.
@@ I Introduction
One appealing property of BDD is that its decoding error probability can be determined analytically [8]: given the number of input errors, the distribution of residual errors after decoding is known. In contrast, no such results exist yet for EaE decoding. In this paper, we derive a closed-form expression for the considered EaED. Specifically, we aim to determine the decoding transition probability (DTP) P(R=r|U=u,E=e){P(R\!=\!r|U\!=\!u,E\!=\!e)}, where $R$ is the number of residual errors after decoding, $U$ the number of input errors, and $E$ the number of input erasures. With the DTP, the post-decoding bit error rate (BER) and frame error rate (FER) for various memoryless channels can be computed. The DTP is validated through simulations and applied in a concatenated scheme.
@@ II Preliminaries
Let $\mathcal{C}[n,k,d_{\textup{des}}]$ be a primitive, narrow-sense, binary BCH code, where $n$ is the block length, $k$ the dimension, and $d_{\textup{des}}$ the designed distance, a lower bound on the minimum distance $d_{\textnormal{min}}$. The code rate is given by $R_{c}=k/n$. Let $t$ denote the number of correctable errors. The parameters are related by $n=2^{b}-1$, $k\geq n-bt$, and $2t+1=d_{\textup{des}}$. Let $A_{w}$ denote the number of codewords of weight $w$ in $\mathcal{C}$. Binary primitive BCH codes have approximately binomial weight distributions given by $A_{w}\approx 2^{-bt}\binom{n}{w}$ for $d_{\mathrm{des}}\leq w\leq n-d_{\mathrm{des}}$, with $A_{0}=A_{n}=1$, and $A_{w}=0$ elsewhere [1]. The exact weight enumerators of some short BCH codes are listed in [9]. Note that both [8] and the results derived in this work rely on the weight distribution. If only an approximate weight distribution is available, the resulting analysis is also approximate.
@@
We use $\bm{x}\in\mathcal{C}$ to denote the transmitted codeword and $\bm{y}$ to denote the received word, whose alphabet is determined by the channel. At the output of a BSC, we have
$\textsf{BDD}\left(\bm{y}\right)=\begin{cases}\bm{c},&\text{if }\bm{y}\in\mathcal{S}_{t}(\bm{c}),\\ \bm{y}\,(\mathsf{fail}),&\text{otherwise},\end{cases}$
where $\mathcal{S}_{t}(\bm{c})=\{\bm{y}\in\mathbb{F}_{2}^{n}:d(\bm{y},\bm{c})\leq t\}$ denote the Hamming sphere of radius $t$ for $\bm{c}\in\mathcal{C}$. The decoding outcome falls into one of three categories: (i) decoding success, if $\bm{c}=\bm{x}$; (ii) decoding failure if no codeword is found; or (iii) miscorrection (also called a decoding error as in [8]), if $\bm{c}\neq\bm{x}$.
@@
Note that for a decoding failure where $\textsf{BDD}\left(\bm{y}\right)=\bm{y}$}, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(R\neq u|u)\!=\!0$}, while for a decoding success, $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(R\neq 0|u)\!=\!0$}. We further denote by $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(u)$, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u)$, and $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(u)$ the conditional probabilities of success, failure, and miscorrection, respectively, given $u$ errors.
@@ III Analysis of Bounded Distance Decoding
We define $a=|\{i\in[n]:c_{i}=1,y_{i}=0\}|$} as the number of correct bits in $\bm{y}$ that are (wrongly) flipped by BDD, and $b=|\{i\in[n]:c_{i}=0,y_{i}=1\}|$} as the number of erroneous positions in $\bm{y}$ that are (correctly) flipped by BDD. Hence, it holds that u+ab=r{u+a-b=r} and $a+b=d\left(\bm{y},\bm{c}\right)\leq t$. For fixed $a$ and $b$, there are $\binom{r}{a}\binom{n-r}{b}=\binom{r}{a}\binom{n-r}{u+a-r}$ distinct patterns for $\bm{y}$. Additionally, the number of codewords $\bm{c}\in\mathcal{C}$ with $\textnormal{wt}\left(\bm{c}\right)=r,$ is given by the weight enumerator $A_{r}$. Thus,
@@ IV-B Decoding Transition Probability
$P_{\textnormal{{EaED}}}^{d}(r|u,e)=\textstyle\sum_{e_{1}=0}^{e}P_{\textnormal{{EaED}}}^{d}(r|u,e,e_{1})\cdot P(e_{1}|u,e),$ (1)
@@
Using the $[255,239,5]$ BCH code as an example, Tables I and II list the conditional probabilities $P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(u,e)$ and $P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(u,e)$, respectively. These quantities represent the probabilities of successful decoding and miscorrection given $u$ errors and $e$ erasures, computed using (1) and marginalized over $R$. Note that if $e=0$, Theorem 2 reverts to Theorem 1. Accordingly, the first column of both tables corresponds to BDD. The miscorrection rate of BDD is high and is close to the upper bound $\frac{1}{t!}$ derived in [8]. Compared with BDD, EaED can decode beyond the minimum distance. When $2u+e$ slightly exceeds $d_{\textnormal{min}}-1$, the decoding success rate of EaED is still high. However, this comes at the expense of an increased probability of miscorrection for larger values of $2u+e$. The post-decoding BERs of EaED and BDD are compared in Fig. 2 with the threshold $\mathsf{T}$ optimized to minimize the BER of EaED at $9$ dB. We see that the simulation and computed values closely agree. Additionally, as expected, EaED improves upon BDD in terms of performance.
@@ V-C Improving one-shot decoding of BCH code
As shown in Fig. 2, EaEDa outperforms both EaED and BDD. Next, we examine the miscorrection rate. For EaED, most bit errors are due to miscorrections, whereas BDD exhibits a smaller fraction. This observation aligns with Tab. I and Tab. II, reflecting that EaED has a stronger error correction capability but is more prone to miscorrection. By contrast, EaEDa substantially reduces miscorrections through the use of anchor bits, which can be seen from the dashed lines in Fig. 2. Thus, EaEDa improves decoding performance while maintaining a low miscorrection rate, keeping the undetectable error rate negligible.
@@ Appendix A EaED Calculation
Finally, we consider the cases in $\mathcal{M}$. First, note that they can only cause a miscorrection or a failure since $\bm{y}^{(1)},\bm{y}^{(2)}\notin\mathcal{S}_{t}({\bm{0}})$. In the following, we make a significant simplification by assuming that the outcomes of $\textsf{BDD}\left(\bm{y}^{(1)}\right)$ and $\textsf{BDD}\left(\bm{y}^{(2)}\right)$, whether miscorrection or failure, are independent of each other. In principle, this assumption is false as $d(\bm{y}^{(1)},\bm{y}^{(2)})=e$. However, as the positions of the erasures are random, we observe that fixing the decoding result of one BDD does not significantly influence the other one. Thus,
@@ Appendix B EaED with Anchor Bits
Acknowledgment
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001899).